# Federated Foundation Models in Harsh Wireless Environments: Prospects, Challenges, and Future Directions

### v1

@@ IV Open Research Directions
Despite the promising capabilities of FFMs in harsh and resource-constrained environments, several key challenges remain unresolved. In this section, we outline critical open problems and propose directions for future research that could further enhance the robustness, adaptability, and effectiveness of FFMs in extreme operational settings. To ground our discussion, we list in Table I several publicly available datasets spanning satellite, forest, and underwater environments that serve as benchmarks for FFMs.
Multimodal Context Alignment: In harsh and resource-constrained environments, FFMs must fuse information from multiple sensing modalities collected across heterogeneous edge devices, such as acoustic, visual, infrared, or environmental signals. In contrast to centralized foundation models, where modalities are well curated and aligned, harsh environments introduce partial modality coverage, severe fidelity variations, and missing modality co-occurrences. For example, identical sensors may degrade differently under high humidity or pressure from harsh environments, producing inconsistent fidelity across clients. Furthermore, local devices may operate with only a subset of the modalities due to limited hardware capabilities or extreme environmental limitations. This partial modality coverage can vary across clients or even over time. As a result, the embedding space can fragment, threatening semantic consistency across the network. Addressing this requires dynamic alignment mechanisms that can stitch together incomplete and noisy cross-modal links, preserve embedding coherence, and adapt to evolving modality availability under strict communication and computation constraints.
Continuous Adaptation in Evolving Conditions: The non-stationarity nature of harsh environments demands continual learning, as underlying data distributions shift due to equipment degradation, structural reconfiguration, or environmental changes. Sensor readings can fluctuate (e.g., pressure gauges drifting over time or optical sensors failing under dust) making catastrophic forgetting even more severe. FFMs must therefore adapt without erasing past knowledge, often in the absence of replay buffers (due to storage limits) or with regularizers that may fail under rapidly changing inputs. Developing robust continual learning strategies that retain knowledge while ensuring long-term adaptability remains a critical research frontier for autonomous FFMs in unpredictable and high-variability settings.
@@
Dataset Environment Modalities Tasks
FLAIR-HUB Satellite and UAV Topographic Information, Satellite Image Time Series(SITS), VHR Land Cover Segmentation, Crop Type Segmentation
Planted Satellite SITS, VHR Forest plantation/species Identification
MADOS Satellite Multispectral Satellite Images Marine debris and oil spill mapping
FinnWoodlands Deep Forest RGB images, point clouds, depth maps Instance Segmentation, Panoptic Segmentation, Depth Completion
SUIM Underwater Image Underwater Species segmentation
LIACI Underwater Image Underwater Ship Inspection
Mitigate Performance Degradation from Drifting Effects: In FFM systems, multiple forms of “drift” can combine to undermine the accuracy and stability of the global model. In FL, data heterogeneity across clients leads to the well‐known “client drift” problem, in which each local model converges toward an optimum that fits its own data distribution, rather than the global optimum.
For FFMs, this challenge becomes even more complex due to the online learning nature of FMs. Two additional drift phenomena emerge: 1) Embedding Drift. The learned embedding gradually changes its semantic meaning over time. This can occur when the underlying data distribution evolves or when the available modality changes, making an embedding learned months ago no longer accurately represent the same concept today. 2) ID Drift. The mapping between discrete identifiers (e.g., item IDs, user IDs) and their corresponding inputs changes over time. In large‐scale systems, such as recommendation or content‐delivery platforms, old IDs frequently expire while new IDs are introduced. As a result, a given embedding vector may be reassigned to represent an entirely different entity, causing severe misalignment between model parameters and the real‐world concepts they are meant to capture. These multi-layered drifts are combined, degrading global performance and stability. Future research must design federated optimization techniques that jointly mitigate heterogeneous drifts while maintaining alignment across distributed, intermittently connected nodes.
Robustness Tuning Under Missing Communication and Corrupted Data: In harsh environments, robustness must be ensured not only against incomplete data but also against unreliable communication and sensor failures. When communication is at its weakest, or in some cases absent altogether, the global model may wait for days to receive a federated update, leaving individual clients isolated. Compounding this issue, sensors are often prone to miscalibration or degradation, producing corrupted, noisy, or even misleading signals that may be unsafe to use for model training. These conditions create severe challenges for FFMs, which must remain resilient in the face of missing updates, extreme data heterogeneity, and corrupted local inputs. Ensuring consistent and safe performance under such conditions remains an open challenge, demanding theoretical guarantees and practical mechanisms to handle both communication breakdowns and unreliable sensing sources.
Energy-Efficient and Sustainable Operation: In harsh environments where power supply is limited or intermittent, ML models must operate within strict energy constraints. Unlike conventional FM training paradigms that primarily evaluate computational cost in terms of floating-point operations per second (FLOPs), local deployment in harsh environments requires careful consideration of total energy consumption over the training and inference lifecycle. To avoid rapid battery depletion on edge devices, a joint optimization of convergence speed and energy usage becomes necessary. In this context, lightweight FM architectures, featuring compact backbones or a minimal number of trainable parameters, are especially advantageous [15]. These models reduce both local computation costs and communication overhead, which is critical given the limited bandwidth and frequent instability of communication links in such environments. Designing energy-aware training schedules and low-power inference strategies will be essential to enable real-time or periodic model updates on resource-constrained platforms, thereby supporting sustained and autonomous operation of FFMs in energy-sensitive deployment scenarios.
@@ V Conclusion
We advocate for the development of Federated Foundation Models (FFMs) for harsh environments as a viable path forward. By combining the generalization power of foundation models with the decentralized, communication-efficient nature of federated learning, FFMs offer a framework for building intelligent systems that are both robust and adaptive in real-world adverse conditions. We have outlined the limitations of conventional methods, highlighted critical system constraints in harsh environments, and presented a vision for FFMs that balances performance, efficiency, and resilience. Looking ahead, progress depends on three directions: (i) co-design with the physical layer, where communication/scheduling protocols are designed with the need of model training; (ii) contact or event-triggered aggregation with hierarchical or semi-decentralized topologies and lightweight adapters/experts for local change without global drift; and (iii) evaluation that goes beyond accuracy to measure energy per improvement, communication per adaptation, recovery after disruption, and stability under continual drift on field datasets. FFMs in harsh environments are thus not merely bigger models distributed farther, they are engineered systems meant to remain dependable when links are unreliable and data are scarce.

### v2

@@ IV Open Research Directions
In this section, we outline critical open problems and propose directions for future research that could further enhance the robustness, adaptability, and effectiveness of FFMs in extreme operational settings.
Multimodal Context Alignment: In harsh environments, sensor noise, modality dropouts, and asynchronous data availability can fragment the embedding space, threatening semantic consistency across the network. We envision environment-aware alignment mechanisms that enable encoders to stitch together incomplete and noisy cross-modal signals, preserving coherence under strict communication and computation constraints. To achieve this, we envision shared latent anchors, a compact set of globally consistent, environment-conditioned reference embeddings that act as semantic landmarks for aligning local representations across clients. These anchors can be periodically broadcast or collaboratively learned, adapting to each client’s sensing and environmental context. To complement this, Cross-Client Contrastive Regularization schemes may allow clients performing related tasks to exchange lightweight prototypes or feature statistics, enabling local contrastive alignment based on prior understanding of client tasks.
@@
[…]
Mitigating Multi-Source Drift: In FFM systems deployed in harsh environments, multiple forms of drift including distributional, embedding, and identifier-level shifts can combine to degrade global model performance and consistency over time. We envision drift-aware FFM optimization frameworks that incorporate continual adaptation modules to track and compensate for evolving semantic representations and ID mappings. For example, embedding replay buffers can be locally maintained to contrast current embeddings with temporally anchored versions, enabling clients to detect and correct semantic drift in their representations. Complementarily, we envision parameter-efficient gradient tracking mechanisms that inject lightweight correction terms into the locally computed stochastic gradients, aligning local updates with globally relevant directions and reducing the effects of complex data distributional drifts across asynchronous clients.
Robustness Tuning Under Missing Communication and Corrupted Data: In harsh environments, robustness must withstand both incomplete data and unreliable communication. Prolonged disconnections can leave clients isolated without updates, while sensor failures or miscalibrations may inject corrupted signals into local training, threatening model reliability. We envision communication-aware robustness tuning mechanisms that adapt learning rates or aggregation weights based on the freshness and reliability of updates, ensuring stable progress and eliminate stale updates even under sparse participation. In parallel, we envision corruption-resilient training objectives that downweight or filter unreliable local signals using noise-aware loss functions or confidence-based data selection, preventing quality degradation of data from effecting the reasoning ability of FFMs.
Energy-Efficient and Sustainable Operation: In harsh environments where power supply is limited or intermittent, FFMs must operate within strict energy budgets for both local computation and communication transmission. We envision energy-aware training schedules that dynamically adjust batch sizes, update frequencies, or participation levels based on real-time device power states (e.g., adapting learning intensity when solar-charged batteries are low). Additionally, we envision client-based FM-driven federation protocol optimization, where the reasoning ability of FMs is leveraged to adapt communication strategies in real time. For example, FFMs could learn semantic communication schemes to compress messages, select efficient channels under acoustic or radio links, or even optimize scheduling and physical-layer operations across the network. Such protocol-level intelligence enables sustainable and self-maintaining federation without constant third-party supervision, ensuring long-term operation under severe energy constraints.
Benchmarking in Realistic Harsh Environments: Current benchmarks [13, 14, 15] (Table I) advance standardized evaluation for federated multimodal systems but lack key features for harsh-environment scenarios. We envision a jointly designed benchmark that incorporates: (i) Environmental stressors (e.g., humidity, corrosion, radiation); (ii) Device-level heterogeneity in compute, memory, and energy constraints; (iii) Robustness testing for partial signal loss and multimodal desynchronization; and (iv) New evaluation protocols with metrics such as energy per update, adaptation latency, and resilience under prolonged disconnection. Such benchmarks are essential to guide the development of FFMs for real-world deployment in extreme settings.
@@ V Conclusion
We advocate for the development of Federated Foundation Models (FFMs) as a promising paradigm for enabling intelligent systems in harsh environments. By combining the generalization capacity of foundation models with the communication-aware nature of federated learning, FFMs offer the potential to support robust, adaptive, and autonomous operation across physically distributed and intermittently connected platforms. Rather than viewing FFMs in harsh environments as simply scaling models to the edge, we argue they must be designed as self-sustaining systems that are capable of learning, adapting, and surviving in the face of unreliability and change.